Day Sample (n) Number (np) Proportion (p)
1 500 12 0.024
2 500 15 0.030
3 500 19 0.038
4 500 13 0.026
5 500 9 0.018
6 500 26 0.052
7 500 18 0.036
8 500 14 0.028
9 500 17 0.034
10 500 18 0.036
11 500 16 0.032
12 500 24 0.048
13 500 11 0.022
14 500 31 0.062
15 500 16 0.032
16 500 10 0.020
17 500 16 0.032
18 500 17 0.034
19 500 20 0.040
20 500 15 0.030
21 500 8 0.016
22 500 13 0.026
23 500 12 0.024
24 500 17 0.034
25 500 18 0.036

Use the collected data above, please construct a p chart and an np chart and explain how do they both look?
Answer #1

For p chart, define p.bar as the mean of the proportions given. Then the control limits for p chart are

UCL=\bar{p}+3\sqrt{\bar{p} (1-\bar{p})/n}\\ CL=\bar{p}\\ LCL=\bar{p}-3\sqrt{\bar{p} (1-\bar{p})/n}\\

We compute p.bar=.0324 and hence

UCL=.05616, CL=.0324 and LCL=.00864

1596089766223_image.png

From the chart, we find that the 14 th observation falls outside the control limits.

For np chart, define p.bar as the mean of the proportions given. Then the control limits for np chart are

UCL=n\bar{p}+3\sqrt{n\bar{p} (1-\bar{p})}\\ CL=n\bar{p}\\ LCL=n\bar{p}-3\sqrt{n\bar{p} (1-\bar{p})}\\

We compute p.bar=.0324 and hence

UCL=28.08, CL=16.2 and LCL=4.32

1596090035373_image.png

From the chart, we find that the 14 th observation also falls outside the control limits.

For any query in above, comment